OpenStudy (anonymous):

a coil of one loop is made by a wire of length L and there after a coil of two loops is made by same wire. the ratio of magnetic field at the centre of coils respectively?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

no ans is 1:4

OpenStudy (anonymous):

twoo loops are placed closed to one another?

OpenStudy (anonymous):

how are two loops placed relative to each other?

OpenStudy (anonymous):

dunno -_-

OpenStudy (anonymous):

ok the formula is uni.....u is meu nd n is coils per unit length nd i is current.........the rest depend on their relative arrangement

OpenStudy (anonymous):

i guess for one loop wire 2pi r = L r=L/2pi b=uoI 2pi r/4pi r^2 B=uo/2r B=uo pi/L B=uo pi/L.................................(1) for 2 loop 4pi r= L r=L/4pi B=2 uoI/2r B=4 uo pi/L.............................(2) from the ratio of eqn 1 and 2 we get 1:4 hurrrah i did it thnx ^.^

OpenStudy (vincent-lyon.fr):

B is proportional to current and inversely proportional to distance (radius). Radius is halved and current is doubled, hence field will be four-fold.